OpenStudy (anonymous):

(Seventh Grade Math - Medal and Fan) 14 is what percent of 56? (Please explain!!)

OpenStudy (anonymous):

25%

jhonyy9 (jhonyy9):

do you know that 56 is 100% yes ?

OpenStudy (anonymous):

@jhonyy9 yes

jhonyy9 (jhonyy9):

so write it in this way using the rule of 3 simple 56 ------ 100% 14 ------ x% ------------------- cross multiplie and calculi the x 56x=14*100 x = 14*100/56 = ? ok. ?

OpenStudy (anonymous):

@jhonyy9 ok ty

jhonyy9 (jhonyy9):

np was my pleasure good luck any time than i m online bye